OpenStudy (anonymous):

Solve 5(2x + 1) = 12x – 1. You must show all work.

Parth (parthkohli):

Distributive property does all the work here.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5(2x) + 5(1) = 12x - 1}\) Then you have a linear equation in a very simple form, which I guess you'd manage.

Parth (parthkohli):

And... Welcome to OpenStudy! http://openstudy.com/code-of-conduct

OpenStudy (mimi_x3):

\[5(2x+1)=12x-1\] \[10x+5=12x-1\] Collect the like terms!! \[10x-12x = -1-5\] The add them!!! then solve for \(x\)

OpenStudy (anonymous):

can you show the full answer? please?

OpenStudy (mimi_x3):

lol i already gave you the full solution. have you tried solving for \(x\)?

OpenStudy (anonymous):

Nah

OpenStudy (mimi_x3):

then do sooo!!

OpenStudy (anonymous):

Ifs for my big brother

OpenStudy (anonymous):

im 10

OpenStudy (mimi_x3):

well, if you want answers you can just use this http://www.wolframalpha.com/ if you want to learn then use this site (:

OpenStudy (anonymous):

god bless your soul

OpenStudy (mimi_x3):

But i advise you to learn instead (:

OpenStudy (anonymous):

thanks you a teacher?

OpenStudy (diyadiya):

Student, But she can teach you :)